A skydiver jumps off of an airplane. The plane flies at an altitude of 4 km, and at a speed of 100 km/hr. The skydiver jumpsNote that the answer is not:

- 0 m/s^2

- 7.848 m/s^2
Answer #1

b> → For the vertical direction Ly-axis), From 4 km to 2.5 km of height, Vy ² - uy² => Vy² -(0) 2 9h (n = 4-2.5 = 1.5 km2 150→ For the horizontal direction (a-anis). From the height of 4 km to 2.5 km the horizontal relocity is 100 km / hr Una = 100 k